OpenStudy (anonymous):

Find the area of a parallelogram with sides of 12 inches and 8 inches if one of the angles is 120

OpenStudy (anonymous):

first calculate the height of the parallelogram sine 60 degrees = height / 8 height = 8 sin60 area = 12 * 8sin 60

OpenStudy (anonymous):

= 12* 8 * sqrt3 /2 = 48sqrt3 = 83.14 sq ins